OpenStudy (anonymous):

can someone help me with 90+3x+(x+10)

OpenStudy (anonymous):

(90+10)=3x + x, (90+10)/3 = 2x, ((90+10)/3), ans/2=x

OpenStudy (gw2011):

Since this problem doesn't ask to find x, I assume it is just to be simplified. Therefore, you get 90+3x+x+10. Combine like terms: 90+10=100 and 3x+x=4x. This then comes out to be: 4x+100